Subtract and simplify. √27 - 5 √12
Could you explain how to subtract and simplify this?
I am not sure if I came up with the correct answer and would like to check my work.
Thank you.
Cynthia

Answer by richwmiller(17219)

actually you have to simplify first
sqrt(27)-5sqrt(12)
sqrt(27)=sqrt(3*3*3)=3sqrt(3)
sqrt(12(=sqrt(2*2*3)=2sqrt(3)
so we have
3sqrt(3)-5*(2sqrt(3)
3sqrt(3)-10sqrt(3)
(3-10)*sqrt(3)
-7sqrt(3)